# LANCAR: Leveraging Language for Context-Aware Robot Locomotion in Unstructured Environments

### v2

@@ Abstract
Navigating robots through unstructured terrains is challenging, primarily due to the dynamic environmental changes. While humans adeptly navigate such terrains by using context from their observations, creating a similar context-aware navigation system for robots is difficult. The essence of the issue lies in the acquisition and interpretation of contextual information, a task complicated by the inherent ambiguity of human language. In this work, we introduce LANCAR, which addresses this issue by combining a context translator with reinforcement learning (RL) agents for context-aware locomotion. LANCAR allows robots to comprehend contextual information through Large Language Models (LLMs) sourced from human observers and convert this information into actionable contextual embeddings. These embeddings, combined with the robot’s sensor data, provide a complete input for the RL agent’s policy network. We provide an extensive evaluation of LANCAR under different levels of contextual ambiguity and compare with alternative methods. The experimental results showcase the superior generalizability and adaptability across different terrains. Notably, LANCAR shows at least a 7.4% increase in episodic reward over the best alternatives, highlighting its potential to enhance robotic navigation in unstructured environments. More details and experiment videos could be found in this link.
@@ I Introduction
Designing locomotion for quadruped robots has been a longstanding focus of research [1]. The variability of environmental physical properties heavily influences the robot’s movement, making it difficult to create a universal policy that works in all situations. Despite significant progress, devising a universal policy capable of effectively addressing all possible scenarios remains elusive  [2, 3]. Previous efforts include using graph-like structures [4] or autoencoders [5] to gather contextual environment information, but these methods often fall short in complex terrain navigation due to limited reasoning capabilities. Addressing these challenges requires innovative approaches, such as combining human insights with technological solutions. Humans can intuitively understand environmental contexts, like associating wet grass with high damping, a concept difficult for current algorithms to grasp. However, leveraging human feedback is complicated by the ambiguity of natural language [6] and the impracticality of expecting humans to provide detailed quantitative descriptions of environments using physical parameters instead of vague qualitative sentences with descriptive words.
@@
Figure 1: Task Description. We consider two robot learning approaches for locomotion guided by ambiguous human descriptions. The first existing approach (TOP) is when the robot moves over diverse terrains with a trained policy without any contextual information. Given the complexity of the terrains, robots may face difficulties in developing a generalized policy to address all types of terrains, leading to the failure of its ultimate policy. Our proposed approach (BOTTOM) has the robot moving over diverse terrains with our trained policy and contextual information from human observers. Robots convert this interpreted contextual information into embeddings with LLM. With the extra contextual information added to robots’ own perceptions from their sensors, robots could develop better policies with a better understanding of the environment.
@@
The recent success of the Large Language Models (LLMs) and their ability to perform chain-of-thought [7], logic reasoning [8], and common sense reasoning [9] is a promising approach to address these problems. Integrating LLMs within Reinforcement Learning (RL) frameworks could enhance RL agents’ learning efficiency and adaptability. RL, critical in robotics for tasks like manipulation [10], navigation [11], and locomotion [12], benefits from interactions with the environment and its adaptability to different scenarios. Techniques like domain randomization [13] have been used to prepare RL agents for diverse conditions. While previous efforts have explored LLMs for predicting RL reward functions [14] or providing robotic control inputs [15], these do not fully leverage LLMs’ reasoning abilities. Instead, positioning LLMs as intermediaries to convert human language into RL-friendly formats could optimize their use, ensuring RL agents’ decision-making is not compromised by direct human instruction.
@@
This study explores the use of LLMs to interpret environmental context, enhancing RL agents’ ability to guide robot locomotion, particularly for a quadruped robot navigating diverse terrains with the assistance of human observers. The context, often unperceivable directly by the robot, includes terrain characteristics crucial for navigation. Fig. 1 gives an overview of our approach. In Scenario $1$, a robot traverses terrains without any contextual information. The result shows the robot is struggling to formulate a universal navigation policy. In Scenario $2$, the robot navigates the same terrains but receives additional contextual information from human observers, [^1] In some cases vision-language models may also be a useful surrogate to provide contextual information that is visual like “You are entering a grassland right after the rain” or “You are walking on a dry rocky road under the sun.” Robots use an LLM-based translator to convert embeddings representing contextual information from human interpretation, enhancing their decision-making process alongside their sensory observations.
@@
Our method, LANCAR (LANguage for Context-Aware Robot locomotion), capitalizes on the versatility of LLMs to interpret human language, transforming it into indices or contextual embeddings. This process mitigates the ambiguity inherent in human language, enabling robots to navigate varied terrains with adaptable, generalized control policies through collaboration with human observers. While this paper focuses on leveraging LLMs instead of Vision Language Models (VLMs) for contextual understanding from images, our approach can readily be extended to incorporate VLMs in practical environments. We validate our approach using the spot-mini-mini robot simulator [16], showing that LANCAR enhances performance compared to a no-context baseline. Specifically, LANCAR with contextual embeddings shows at least 7.3% improvement on episodic rewards in low-level tasks and 7.5% improvement in high-level tasks over the baselines. Additionally, LANCAR using the indexing feature leads to a 1.7% improvement in high-level tasks compared to the no-context approach.
@@
  • We propose a novel approach, LANCAR, that incorporates LLMs into RL in robot decision-making that enables robots to understand external contextual information from human observers and generate a more robust and generalized RL policy.
  • We propose an LLM-based contextual information translator module that interprets high-level, ambiguous, human language contextual information of environments into contextual information embeddings accessible for RL agents with the reasoning ability of LLMs.
  • We evaluate LANCAR with four different backbone RL approaches under $10$ case studies using both low-level and high-level contextual information. We validate the efficacy of LANCAR in policy generalizability and adaptability across diverse terrains which shows at least 7.3% and 7.5% of performance improvement over established baselines and over $10$ times higher episodic reward than ablations using different backbone RL approaches.
@@ III-A Problem Formation
We model the problem as an extension of a Partially Observable Markov Decision Process (POMDP), specifically as an implicit POMDP [52]. An implicit POMDP is specified by a tuple, $\left\langle\mathcal{S},\mathcal{A},\mathcal{O},\Omega,\mathcal{Z},\mathcal{F},\mathcal{T},\mathcal{R},\gamma\right\rangle$, where the state space, $\mathcal{S}=\mathcal{S}_{ex}\cup\mathcal{S}_{im}$, is composed of both explicitly observable states $\mathcal{S}_{ex}$ and implicitly observable states $\mathcal{S}_{im}$. The explicitly observable states are those environmental states directly observable from the agent’s onboard sensors. The agent’s observation space is $\mathcal{O}$. The observation function is given by $\Omega:\mathcal{S}_{ex}\rightarrow\mathcal{O}$. The implicitly observable states are the contextual information in the environment that cannot be detected directly by the robots but still affect robots’ policies. $\mathcal{Z}$ denotes the embedding of the contextual information from the implicitly observable states $\mathcal{S}_{im}$, while the mapping function between the two is $\mathcal{F}:\mathcal{S}_{im}\rightarrow\mathcal{Z}$. Nevertheless, the implicitly observable states (i.e. contextual information) can still be inferred by robots through reasoning over visual perception or tactile sensing or through human language feedback. In this work, our primary focus is to recover $\mathcal{S}_{im}$ using contextual information given in natural language.
@@
Figure 2: Context-Aware Reinforcement Learning Robot Locomotion. Our framework adds a context translator to the RL setup, enhancing navigation across varied terrains. Agents receive direct observations from the environment, while human observers provide context information, interpreting terrain properties into natural language. The LLM translator processes this into contextual embeddings, merged with direct observations for RL agent input. The agents then apply their control policies to these enriched inputs to determine and perform actions in the environment.
@@ III-B Human-Robot Collaboration Framework
We introduce a human-robot collaboration framework, LANCAR, as depicted in Figure 2. To recover the contextual information from environments, we introduce the LLM-based context translator module in addition to the standard RL agent. When robots traverse in environments with diverse terrains at time $t$, robots observe the environment’s explicitly observable states $s_{ex}^{t}$, and the human observer interprets the implicitly observable states $s_{im}$ (i.e. contextual information). Here, we assume that the contextual information is consistent within one episode so that $s_{im}$ is fixed. The human observer provides qualitative descriptions or captions of the contextual information to the LLM translator. The LLM translator extracts the environmental properties from the contextual information and generates the contextual embedding $z$, which is concatenated with the observations $o_{t}$ as the input for RL agents. RL agents produce the action $a_{t}$ using their control policies $\pi$ given the context-aware inputs and execute the action in the environment. This framework is designed to be compatible with the other RL methods, offering flexibility in its implementation.
@@ III-C LLM-based Context Translator
In our framework, the LLM is pivotal, converting human-interpreted environmental contextual information into embeddings that RL agents can directly use. We crafted a context translator module leveraging In-context Learning (ICL), enabling the LLM to use its reasoning capabilities through zero-shot or few-shot examples, thus facilitating an interpretable way to interact with the LLM sans explicit training, akin to mimicking human reasoning and decision-making processes [40].
We feed the LLM descriptive sentences about the environment’s contextual information and accompany these with prompts that include examples of potential inputs and outputs the model encounter. An example prompt is presented in Figure 3. These prompts are structured to guide the LLM in mapping qualitative environmental descriptions to embeddings through a series of multiple-choice questions. Each question pertains to a specific environmental characteristic, with the LLM tasked with selecting from pre-defined qualitative descriptors. The chosen answers are then transformed into concatenated one-hot vectors, creating the contextual embeddings for RL agents.
By providing in-context examples, we aid the LLM in grasping the task’s nature. Upon receiving inputs, the LLM is expected to respond to the questions based on the established format of the in-context examples, ensuring the generation of relevant contextual embeddings.
Let $\oplus:X\times Y\rightarrow[X,Y]$ be an operator concatenating two vectors, and let $\text{onehot}(x)$ denote the function for one-hot encoding. For each property $i$, it could have different levels in a set $P_{i}$ which include Very Low, Low, Medium, High and Very High. Given $n$ properties and their corresponding levels $v_{p_{i}}$ for property $i$, the contextual embedding of this terrain, denoted as $C$, is obtained by considering the concatenation of the one-hot encoding of the property indexes:
@@
For example, if the contextual information describes the terrain with two properties, saying This terrain has very low friction and very high damping. very low friction maps into an one-hot vector $[1,0,0,0,0]$ and very low friction maps into another one-hot vector $[0,0,0,0,1]$, then the contextual embedding of this terrain is $[1,0,0,0,0,0,0,0,0,1]$.
We note that our contextual approach, leveraging human-generated prompts and responses, enables the LLM to effectively bridge the gap between natural language descriptions and actionable state information, a key aspect of our framework’s success in recovering contextual information from unobservable states of environments.
Figure 3: An Example Prompt for LANCAR. The prompt for LANCAR consists of five sections. The first section outlines the high-level task for the LLM. The second provides details and examples of relevant terrain properties. The third includes in-context learning examples, featuring low-level terrain contexts with outputs derived from multiple-choice question-answering. The final two sections involve presenting inputs to the LLM to generate contextual embeddings and the corresponding outputs.
@@ IV Empirical Results and Discussion
In the experiments, we aim to answer the following two questions regarding performance and policy generalization. The first question is: Does external contextual information improve the performance of the agent when the agent is operating in diverse conditions? To answer this question, we designed a series of experiments and compared our framework with alternative approaches that include or exclude contextual information. In these experiments, each context is a different terrain. The second question we investigate is: Is the LLM model effective when the given input is high-level, open-ended, and ambiguous in retrieving contextual information and thereby in robot locomotion?. To answer this question, we use low-level, precise, and organized human interpretation of contextual information in training but also use high-level, vague, and unorganized contextual information from human observers in evaluation, apart from the low-level contextual information evaluation cases.
We use GPT-4 [57] as our LLM model to interpret both low-level and high-level human instructions for robots into robot-understandable embeddings with a series of formulated multiple-choice questions. All reinforcement learning agents are trained under domain randomization manner with $8$ different scenarios, all of which have diverse environmental properties. The episodic reward curve during training is present in Figure. 4. After the training process, agents are evaluated under $10$ evaluation cases with $5$ cases using low-level contextual information from the environment and $5$ cases using high-level contextual information. All evaluation results are averaged over $16$ episodes.
@@ IV-A Environments
We use a quadruped robot locomotion simulator, spot-mini-mini [16], built in PyBullet [58]. The robot’s goal is to advance along the x-axis as much as possible within a fixed set timeframe of 5,000 steps, minimizing deviation from this axis. The raw observation state from the environment is $16$–dimensional, including the robot’s roll, pitch, gyroscopes, and acceleration, as well as a $4$-length binary vector denoting the robot’s leg contacts with the ground. The extra observation state, depending on the embedding mechanism covered in Section IV-D, is also provided to the agent. The action space is the desired joint angle for each of the $14$ joints that are clipped within the maximum allowed velocities. The reward function is given by the combination of multiple terms corresponding to the robot’s state, including the robot’s traveling distance $d_{x}$, denoting the accomplished distance from the origin in $x$-axis, the penalty $d_{y}$ that represents the robot’s deviation from the $y$ axis, and the penalty $r_{p}$ that happens when the robot does not keep the desired rate. The reward function is defined as $J=d_{x}+0.03d_{y}+10r_{p}$. For the ARS agent using in LANCAR, the learning rate is $0.03$. The number of samples for $\delta$ is $16$. The noise amplitude applied in the exploration is $0.05$.
@@
Manually crafting contextual information for each training instance is unfeasible due to the vast amount of data. Instead, we employ an LLM to automate the generation of detailed, low-level terrain descriptions. This process starts with generating random samples of the terrain’s parameters quantitatively describing properties given in Table I, followed by a prompt instructing the LLM to translate these values into qualitative descriptions—ranging from Very Low to Very High, based on $18$ in-context learning examples. Given the actual parameter values, we ask the LLM to generate a low-level contextual description. A sample description generated during training is: This environment has very low restitution when collision, low friction, very high stiffness level, and very low damping.

### v3

@@ Abstract
Navigating robots through unstructured terrains is challenging, primarily due to the dynamic environmental changes. While humans adeptly navigate such terrains by using context from their observations, creating a similar context-aware navigation system for robots is difficult. The essence of the issue lies in the acquisition and interpretation of context information, a task complicated by the inherent ambiguity of human language. In this work, we introduce LANCAR, which addresses this issue by combining a context translator with reinforcement learning (RL) agents for context-aware locomotion. LANCAR allows robots to comprehend context information through Large Language Models (LLMs) sourced from human observers and convert this information into actionable context embeddings. These embeddings, combined with the robot’s sensor data, provide a complete input for the RL agent’s policy network. We provide an extensive evaluation of LANCAR under different levels of context ambiguity and compare with alternative methods. The experimental results showcase the superior generalizability and adaptability across different terrains. Notably, LANCAR shows at least a 7.4% increase in episodic reward over the best alternatives, highlighting its potential to enhance robotic navigation in unstructured environments. More details and experiment videos could be found in this link.
@@ I Introduction
Designing locomotion for quadruped robots has been a longstanding focus of research [1]. The variability of environmental physical properties heavily influences the robot’s movement, making it difficult to create a universal policy that works in all situations. Despite significant progress, devising a universal policy capable of effectively addressing all possible scenarios remains elusive  [2, 3]. Previous efforts include using graph-like structures [4] or autoencoders [5] to gather context environment information, but these methods often fall short in complex terrain navigation due to limited reasoning capabilities. Addressing these challenges requires innovative approaches, such as combining human insights with technological solutions. Humans can intuitively understand environmental contexts, like associating wet grass with high damping, a concept difficult for current algorithms to grasp. However, leveraging human feedback is complicated by the ambiguity of natural language [6] and the impracticality of expecting humans to provide detailed quantitative descriptions of environments using physical parameters instead of vague qualitative sentences with descriptive words.
@@
Figure 1: Task Description. We consider two robot learning approaches for locomotion guided by ambiguous human descriptions. The first existing approach (TOP) is when the robot moves over diverse terrains with a trained policy without any context information. Given the complexity of the terrains, robots may face difficulties in developing a generalized policy to address all types of terrains, leading to the failure of its ultimate policy. Our proposed approach (BOTTOM) has the robot moving over diverse terrains with our trained policy and context information from human observers. Robots convert this interpreted context information into embeddings with LLM. With the extra context information added to robots’ own perceptions from their sensors, robots could develop better policies with a better understanding of the environment.
@@
The recent success of Large Language Models (LLMs) and their ability to perform chain-of-thought [7], logical reasoning [8], and common sense reasoning [9] offers a promising solution to these challenges. Techniques such as domain randomization [10] have been applied to prepare RL agents for varied conditions. Although prior work has explored using LLMs for predicting RL reward functions [11] or providing robotic control inputs [12], these approaches do not fully utilize LLMs’ reasoning capabilities. Positioning LLMs as intermediaries to translate human language into RL-compatible formats could optimize their effectiveness, preventing RL agents’ decision-making from being directly influenced by human instructions.
@@
This study explores the use of LLMs to interpret environmental context, enhancing RL agents’ ability to guide robot locomotion, particularly for a quadruped robot navigating diverse terrains with the assistance of human observers. The context, often unperceivable directly by the robot, includes terrain characteristics crucial for navigation. Fig. 1 gives an overview of our approach. In Scenario $1$, a robot traverses terrains without any context information. The result shows the robot is struggling to formulate a universal navigation policy. In Scenario $2$, the robot navigates the same terrains but receives additional context information from human observers, [^1] In some cases vision-language models may also be a useful surrogate to provide context information that is visual like “You are entering a grassland right after the rain” or “You are walking on a dry rocky road under the sun.” Robots use an LLM-based translator to convert embeddings representing context information from human interpretation, enhancing their decision-making process alongside their sensory observations.
@@
Our method, LANCAR (LANguage for Context-Aware Robot locomotion), capitalizes on the versatility of LLMs to interpret human language, transforming it into indices or context embeddings. This process mitigates the ambiguity inherent in human language, enabling robots to navigate varied terrains with adaptable, generalized control policies through collaboration with human observers. While this paper focuses on leveraging LLMs instead of Vision Language Models (VLMs) for context understanding from images, our approach can readily be extended to incorporate VLMs in practical environments. We validate our approach using the spot-mini-mini robot simulator v.2.1.0 [13], showing that LANCAR enhances performance compared to a no-context baseline. Specifically, LANCAR with context embeddings shows at least 7.3% improvement on episodic rewards in low-level tasks and 7.5% improvement in high-level tasks over the baselines. Additionally, LANCAR using the indexing feature leads to a 1.7% improvement in high-level tasks compared to the no-context approach.
@@
  • We propose a novel approach, LANCAR, that incorporates LLMs into RL in robot decision-making that enables robots to understand external context information from human observers and generate a more robust and generalized RL policy.
  • We propose an LLM-based context information translator module that interprets high-level, ambiguous, human language context information of environments into context information embeddings accessible for RL agents with the reasoning ability of LLMs.
  • We evaluate LANCAR with four different backbone RL approaches under $10$ case studies using both low-level and high-level context information. We validate the efficacy of LANCAR in policy generalizability and adaptability across diverse terrains which shows at least 7.3% and 7.5% of performance improvement over established baselines and over $10$ times higher episodic reward than ablations using different backbone RL approaches.
@@ III-A Problem Formation
We model the problem as an extension of a Partially Observable Markov Decision Process (POMDP), specifically as an implicit POMDP [50]. An implicit POMDP is specified by a tuple, $\left\langle\mathcal{S},\mathcal{A},\mathcal{O},\Omega,\mathcal{Z},\mathcal{F},\mathcal{T},\mathcal{R},\gamma\right\rangle$, where the state space, $\mathcal{S}=\mathcal{S}_{ex}\cup\mathcal{S}_{im}$, is composed of both explicitly observable states $\mathcal{S}_{ex}$ and implicitly observable states $\mathcal{S}_{im}$. The explicitly observable states are those environmental states directly observable from the agent’s onboard sensors. The agent’s observation space is $\mathcal{O}$. The observation function is given by $\Omega:\mathcal{S}_{ex}\rightarrow\mathcal{O}$. The implicitly observable states are the context information in the environment that cannot be detected directly by the robots but still affect robots’ policies. $\mathcal{Z}$ denotes the embedding of the context information from the implicitly observable states $\mathcal{S}_{im}$, while the mapping function between the two is $\mathcal{F}:\mathcal{S}_{im}\rightarrow\mathcal{Z}$. Nevertheless, the implicitly observable states (i.e. context information) can still be inferred by robots through reasoning over visual perception or tactile sensing or through human language feedback. In this work, our primary focus is to recover $\mathcal{S}_{im}$ using context information given in natural language.
@@
Figure 2: Context-Aware Reinforcement Learning Robot Locomotion. Our framework adds a context translator to the RL setup, enhancing navigation across varied terrains. Agents receive direct observations from the environment, while human observers provide context information, interpreting terrain properties into natural language. The LLM translator processes this into context embeddings, merged with direct observations for RL agent input. The agents then apply their control policies to these enriched inputs to determine and perform actions in the environment.
@@ III-B Human-Robot Collaboration Framework
We introduce a human-robot collaboration framework, LANCAR, as depicted in Fig. 2. To recover the context information from environments, we introduce the LLM-based context translator module in addition to the standard RL agent. When robots traverse in environments with diverse terrains at time $t$, robots observe the environment’s explicitly observable states $s_{ex}^{t}$, and the human observer interprets the implicitly observable states $s_{im}$ (i.e. context information). Here, we assume that the context information is consistent within one episode so that $s_{im}$ is fixed. The human observer provides qualitative descriptions or captions of the context information to the LLM translator. The LLM translator extracts the environmental properties from the context information and generates the context embedding $z$, which is concatenated with the observations $o_{t}$ as the input for RL agents. RL agents produce the action $a_{t}$ using their control policies $\pi$ given the context-aware inputs and execute the action in the environment. This framework is designed to be compatible with the other RL methods, offering flexibility in its implementation.
@@ III-C LLM-based Context Translator
In our framework, the LLM is pivotal, converting human-interpreted environmental context information into embeddings that RL agents can directly use. We crafted a context translator module leveraging In-context Learning (ICL), enabling the LLM to use its reasoning capabilities through zero-shot or few-shot examples, thus facilitating an interpretable way to interact with the LLM sans explicit training, akin to mimicking human reasoning and decision-making processes [38].
We feed the LLM descriptive sentences about the environment’s context information and accompany these with prompts that include examples of potential inputs and outputs the model encounter. An example prompt is presented in Fig. 3. These prompts are structured to guide the LLM in mapping qualitative environmental descriptions to embeddings through a series of multiple-choice questions. Each question pertains to a specific environmental characteristic, with the LLM tasked with selecting from pre-defined qualitative descriptors. The chosen answers are then transformed into concatenated one-hot vectors, creating the context embeddings for RL agents.
By providing in-context examples, we aid the LLM in grasping the task’s nature. Upon receiving inputs, the LLM is expected to respond to the questions based on the established format of the in-context examples, ensuring the generation of relevant context embeddings.
Let $\oplus:X\times Y\rightarrow[X,Y]$ be an operator concatenating two vectors, and let $\text{onehot}(x)$ denote the function for one-hot encoding. For each property $i$, it could have different levels in a set $P_{i}$ which include Very Low, Low, Medium, High and Very High. Given $n$ properties and their corresponding levels $v_{p_{i}}$ for property $i$, the context embedding of this terrain, denoted as $C$, is obtained by considering the concatenation of the one-hot encoding of the property indexes:
@@
For example, if the context information describes the terrain with two properties, saying This terrain has very low friction and very high damping. Very low friction maps into an one-hot vector $[1,0,0,0,0]$ and very high damping maps into another one-hot vector $[0,0,0,0,1]$, then the context embedding of this terrain is $[1,0,0,0,0,0,0,0,0,1]$.
We note that our context approach, leveraging human-generated prompts and responses, enables the LLM to effectively bridge the gap between natural language descriptions and actionable state information, a key aspect of our framework’s success in recovering context information from unobservable states of environments.
Figure 3: An Example Prompt for LANCAR. The prompt for LANCAR consists of five sections. The first section outlines the high-level task for the LLM. The second provides details and examples of relevant terrain properties. The third includes in-context learning examples, featuring low-level terrain contexts with outputs derived from multiple-choice question-answering. The final two sections involve presenting inputs to the LLM to generate context embeddings and the corresponding outputs.
@@ IV Empirical Results and Discussion
In the experiments, we aim to answer the following two questions regarding performance and policy generalization. The first question is: Does external context information improve the performance of the agent when the agent is operating in diverse conditions? To answer this question, we designed a series of experiments and compared our framework with alternative approaches that include or exclude context information. In these experiments, each context is a different terrain. The second question we investigate is: Is the LLM model effective when the given input is high-level, open-ended, and ambiguous in retrieving context information and thereby in robot locomotion?. To answer this question, we use low-level, precise, and organized human interpretation of context information in training but also use high-level, vague, and unorganized context information from human observers in evaluation, apart from the low-level context information evaluation cases.
We use GPT-4 [55] as our LLM model to interpret both low-level and high-level human instructions for robots into robot-understandable embeddings with a series of formulated multiple-choice questions. All reinforcement learning agents are trained under domain randomization manner with $8$ different scenarios, all of which have diverse environmental properties. The episodic reward curve during training is present in Fig. 4. After the training process, agents are evaluated under $10$ evaluation cases with $5$ cases using low-level context information from the environment and $5$ cases using high-level context information. All evaluation results are averaged over $16$ episodes.
@@ IV-A Environments
We use a quadruped robot locomotion simulator, spot-mini-mini v.2.1.0 [13], built in PyBullet [56]. The robot’s goal is to advance along the x-axis as much as possible within a fixed set timeframe of 5,000 steps, minimizing deviation from this axis. The raw observation state from the environment is $16$–dimensional, including the robot’s roll, pitch, gyroscopes, and acceleration, as well as a $4$-length binary vector denoting the robot’s leg contacts with the ground. The extra observation state, depending on the embedding mechanism covered in Section IV-D, is also provided to the agent. The action space is the desired joint angle for each of the $14$ joints that are clipped within the maximum allowed velocities. The reward function is given by the combination of multiple terms corresponding to the robot’s state, including the robot’s traveling distance $d_{x}$, denoting the accomplished distance from the origin in $x$-axis, the penalty $d_{y}$ that represents the robot’s deviation from the $y$ axis, and the penalty $r_{p}$ that happens when the robot does not keep the desired rate. The reward function is defined as $J=d_{x}+0.03d_{y}+10r_{p}$. For the ARS agent using in LANCAR, the learning rate is $0.03$. The number of samples for $\delta$ is $16$. The noise amplitude applied in the exploration is $0.05$.
@@
Manually crafting context information for each training instance is unfeasible due to the vast amount of data. Instead, we employ an LLM to automate the generation of detailed, low-level terrain descriptions. This process starts with generating random samples of the terrain’s parameters quantitatively describing properties given in Table I, followed by a prompt instructing the LLM to translate these values into qualitative descriptions—ranging from Very Low to Very High, based on $18$ in-context learning examples. Given the actual parameter values, we ask the LLM to generate a low-level context description. A sample description generated during training is: This environment has very low restitution when collision, low friction, very high stiffness level, and very low damping.